D1 - Back in Position M.Ellis - VC124 - 27th August 2009 Willie & team had shield completely removed and upstream beam line restored in less than four days!!!!! 3
Upstream Beamline - Tested M.Ellis - VC124 - 27th August 2009 All lights work. All PSUs work. Magnets run at large current for several hours with no problems 4
Decay Solenoid - Tested M.Ellis - VC124 - 27th August 2009 48 Hour Soak Test Full Current (870 A) No Problems! Additional checks showed good margin. see MICE News and elog for more. 5
Downstream Beamline - Tested 5 hour soak test. Also run with decay solenoid and upstream magnets yesterday. Remaining niggle – PSUs occasionally trip due to lack of water – investigations under way.
Target 1 – In ISIS M.Ellis - VC124 - 27th August 20097
Target 2 – In R78 M.Ellis - VC124 - 27th August 2009 See Target Talk for Full Details.
Issues / Problems Work has mostly gone very smoothly. My main concern is over communication. We are shortly returning to a period of data taking and it has become clear that it is not always easy to find the expert you need when you need them. It has also been the case that work has been done which should have involved wider consultation (due to the impact on other groups). This is NOT a call for more meetings, but rather a plea for people to keep everyone “in the loop” as to what they are doing and most importantly, what they plan to do. M.Ellis - VC124 - 27th August 20099
Beamline Phone Based on a successful model used by the tracker team. A team of people are being trained up to be expert on the complete beam line (target & all magnets). This team will take it in turns to be on call with a dedicated phone. The MOM and those on shift will always know that they can call one number and get assistance with whatever they need relating to the beam line. M.Ellis - VC124 - 27th August 200910
Next Steps Powering up GVA1 and TOF0 to ensure they are ready for data taking. Test of GVA1 with the decay solenoid powered up (small concern about field near one end of the PMT). Data mover work (Brunel summer student) is progressing nicely. Clean up of DSA. Floor work (epoxying) is ongoing. Complete the commissioning and testing of target 1 (in ISIS) and running of target 2 (in R78). Finish (and test!) documentation for upcoming runs. Data taking under the leadership of the next MOM (Linda Coney) with the assistance of Terry (TTT) Tate. M.Ellis - VC124 - 27th August 200911